### v1

@@ 1 Introduction
In the spirit of previous work [HKJ+22, JGS23, YLS+24, MPRT24], we model SEO as a game and define a natural class of ranking functions we term proportional ranking functions. While in our model a ranking function is a vector-valued function (as it assigns exposure probabilities for each publisher), a proportional ranking function is uniquely defined by a one-variable scalar function, which we call an activation function. Importantly, the choice of the ranking functions determines the incentives of the strategic publishers and hence affects the content available to users in the corpus. Relying on the notion of socially concave games, a subclass of concave games presented in [EdMN09], we prove that the concavity of the activation function guarantees the convergence of any no-regret learning dynamics in any induced SEO game. To do so, we establish a rather stronger statement by proving the equivalence of the following three conditions: the activation function is concave; any induced game is socially-concave; any induced game is concave (as in [Ros65]). We also study the empirical trade-offs between users’ and publishers’ welfare under different choices of the activation function, using a state-of-the-art no-regret dynamics algorithm [FAL+22]. In addition, we demonstrate the effect of the game parameters such as the penalty factor and the number of publishers on the publishers’ welfare, the users’ welfare and the convergence rate.
@@ 2 Preliminaries
[EdMN09] showed that every socially-concave game is a concave game, but the reverse direction is not necessarily true. [EdMN09] also showed that if each player follows any regret minimization procedure in a socially-concave game, then the dynamics are guaranteed to converge. The convergence is in the sense that both the average action vector converges to a Nash equilibrium and the utility of each player converges to her utility in that Nash equilibrium.
@@ 3 The model
Each publisher $i$ tries to achieve two goals simultaneously: on the one hand, maximize the expected chance of her being ranked first by the search engine, which for a given strategy profile $x$ is $\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{r_{i}(x;x^{*})}$; and on the other hand, provide content $x_{i}$ which is as faithful as possible to her initial document, meaning with $d(x_{i},x_{0}^{i})$ as small as possible. Her trade-off between these two goals is determined by an integrity parameter $\lambda_{i}>0$.
@@
$u_{i}(x)\coloneqq\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{r_{i}(x;x^{*})}-\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ (2)
@@
Note that condition A2. of social concavity holds with $\alpha_{i}=\frac{1}{n}$ in every publishers’ game, since $\sum_{i=1}^{n}r_{i}\equiv 1$. This leads to the following useful lemma, whose proof is deferred to Appendix A.1.
@@ 4 Main result
Theorem 1, together with the findings of [EdMN09] that in socially-concave games any no-regret dynamics converges, and the guarantee of [Ros65] that concave games possess a NE and strictly concave games possess a unique NE, leads to the following result:
@@ 5.1 Simulation details
$\mathcal{V}(x)\coloneqq\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{\sum_{i=1}^{n}\bigl{(}1-d(x_{i},x^{*})\bigr{)}\cdot r_{i}(x)}=1-\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{\sum_{i=1}^{n}d(x_{i},x^{*})\cdot r_{i}(x)}$ (5)
@@ 5.2.2 Comparison within ranking function families
Beyond the comparison between the three ranking function families, tuning the hyperparameters within each family also reveals interesting phenomena. The experimental results supporting these observations are provided in Appendix B, and are discussed here. The results once again underscore the publisher-user trade-off discussed in §5.2.1: hyperparameter values that lead to higher publishers’ welfare result in lower users’ welfare and slower convergence rate, and vice versa. As was the case in the comparison between ranking function families, the hyperparameters’ values that lead to higher publishers’ welfare (and lower users’ welfare) are the values that make the ranking function closer to uniform ranking. For example, a decrease in the power parameter $a$ of $r^{root}$ can be seen in Figure 6 to correlate with an increase in publishers’ and a decrease in users’ welfare, and indeed when the power parameter $a$ approaches 00, the root proportional ranking function approaches a uniform ranking. That is: $r^{lin}_{i}\to\frac{1}{n}$ as $a\to 0$. Similarly, both the intercept parameter $b$ of $r^{lin}$ and the shift parameter $c$ of $r^{log}$ are positively correlated with the publishers’ welfare and negatively correlated with the users’ welfare, and indeed $r^{lin}_{i}\to\frac{1}{n}$ as $b\to\infty$ and $r^{log}_{i}\to\frac{1}{n}$ as $c\to\infty$.
@@ 6 Discussion
We studied an information retrieval game in which publishers have initial documents they prefer to provide and initiated the analysis of no-regret dynamics in such games. We defined the class of proportional ranking functions, which are ranking functions that are determined by a one-variable scalar activation function, and established a full characterization of which proportional ranking functions induce socially-concave games and which induce concave games. We then used this characterization to conclude that any concave activation function guarantees the convergence of no-regret dynamics, thereby offering a system designer a rich family of ranking functions that, under the assumption that publishers minimize their regret, guarantee the reach of the system to a stable state. We then empirically investigated the publishers’ welfare and the users’ welfare, as well as the converge rate, which measures how long it takes the ecosystem to reach stability. In particular, we examined how these three criteria are affected by the choice of the ranking function and by changes in the ecosystem structure.
@@ A.1 Proof of Lemma 1
$g(x)$ $\coloneqq\sum_{i=1}^{n}\alpha_{i}\cdot u_{i}(x)=\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{\sum_{i=1}^{n}\frac{1}{n}\cdot\left(r_{i}(x;x^{*})-\lambda_{i}\cdot d(x_{i},x_{0}^{i})\right)}$ (6)
@@
since $\sum_{i=1}^{n}r_{i}(x;x^{*})=1$ for any strategy profile $x\in X$ and information need $x^{*}\in[0,1]^{k}$. Now, $g(x)$ is concave since for every $i\in N$ and $x_{0}^{i}\in[0,1]^{k}$, $d(x_{i},x_{0}^{i})$ is convex in $x_{i}$. Regarding condition A2., it can be easily shown (using the linearity and monotonicity of expectation) that the convexity of $r_{i}(x;x^{*})$ in $x_{-i}$ for any $x_{i}\in X_{i},x^{*}\in[0,1]^{k}$ implies that $\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{r_{i}(x;x^{*})}$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. Together with the fact that $d(x_{i},x_{0}^{i})$ is fixed in $x_{-i}$, we get that $u_{i}(x)=\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{r_{i}(x;x^{*})-\lambda_{i}\cdot d(x_{i},x_{0}^{i})}$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. ∎
A.2 Full derivation of inequality $f^{\prime\prime}(\hat{a})>0$ from the main proof
This appendix provides a detailed derivation of the equation $f^{\prime\prime}(\hat{a})>0$ from the last lines of the main proof, Proof 4. By the definition of the utility in publishers’ games, and since in the game we constructed the distribution over the information needs is a one-hot distribution that places probability $1$ on $x^{*}=0$:
@@ A.3 Proof that if $d$ is strictly bi-convex and $g$ is concave then $r$ induces strictly concave games
$u_{i}(x_{i},x_{-i})\coloneqq\operatorname*{\mathbb{E}}[x^{*}\sim P^{*}]{r_{i}(x;x^{*})}-\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ (18)

### v2

@@ 1 Introduction
In the spirit of previous work [HKJ+22, JGS23, YLS+24, MPRT24], we model SEO as a game and define a natural class of ranking functions we term proportional ranking functions. While in our model a ranking function is a vector-valued function (as it assigns exposure probabilities for each publisher), a proportional ranking function is uniquely defined by a one-variable scalar function, which we call an activation function. Importantly, the choice of the ranking function determines the incentives of the strategic publishers and hence affects the content available to users in the corpus. Relying on the notion of socially concave games, a subclass of concave games presented in [EdMN09], we prove that the concavity of the activation function guarantees the convergence of any no-regret learning dynamics in any induced SEO game. To do so, we establish a rather stronger statement by proving the equivalence of the following three conditions: the activation function is concave; any induced game is socially-concave; any induced game is concave (as in [Ros65]). We also study the empirical trade-offs between users’ and publishers’ welfare under different choices of the activation function, using a state-of-the-art no-regret dynamics algorithm [FAL+22]. In addition, we demonstrate the effect of the game parameters such as the penalty factor and the number of publishers on the publishers’ welfare, the users’ welfare and the convergence rate.
@@ 2 Preliminaries
[EdMN09] showed that every socially-concave game is a concave game, but the reverse direction is not necessarily true. [EdMN09] also showed that if each player follows any regret minimization procedure in a socially-concave game, then the dynamics are guaranteed to converge. The convergence is in the sense that both the average action vector converges to a Nash equilibrium and the average utility of each player converges to her utility in that Nash equilibrium.
@@ 3 The model
Each publisher $i$ tries to achieve two goals simultaneously: on the one hand, maximize the expected chance of her being ranked first by the search engine, which for a given strategy profile $x$ is $\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]$; and on the other hand, provide content $x_{i}$ which is as faithful as possible to her initial document, meaning with $d(x_{i},x_{0}^{i})$ as small as possible. Her trade-off between these two goals is determined by an integrity parameter $\lambda_{i}>0$.
@@
$u_{i}(x)\coloneqq\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]-\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ (2)
@@
Note that condition A1. of social concavity holds with $\alpha_{i}=\frac{1}{n}$ in every publishers’ game, since $\sum_{i=1}^{n}r_{i}\equiv 1$. This leads to the following useful lemma, whose proof is deferred to Appendix A.1.
@@ 4 Main result
Theorem 1, together with the findings of [EdMN09] that in socially-concave games any no-regret dynamics converges, and the guarantee of [Ros65] that (strictly) concave games possess a (unique) NE, leads to the following result:
@@ 5.1 Simulation details
$\mathcal{V}(x)\coloneqq\mathbb{E}_{x^{*}\sim P^{*}}\left[\sum_{i=1}^{n}\bigl{(}1-d(x_{i},x^{*})\bigr{)}\cdot r_{i}(x)\right]=1-\mathbb{E}_{x^{*}\sim P^{*}}\left[\sum_{i=1}^{n}d(x_{i},x^{*})\cdot r_{i}(x)\right]$ (5)
@@ 5.2.2 Comparison within ranking function families
Beyond the comparison between the three ranking function families, tuning the hyperparameters within each family also reveals interesting phenomena. The experimental results supporting these observations are provided in Appendix B, and are discussed here. The results once again underscore the publisher-user trade-off discussed in §5.2.1: hyperparameter values that lead to higher publishers’ welfare result in lower users’ welfare and slower convergence rate, and vice versa. As was the case in the comparison between ranking function families, the hyperparameters’ values that lead to higher publishers’ welfare (and lower users’ welfare) are the values that make the ranking function closer to uniform ranking. For example, a decrease in the power parameter $a$ of $r^{root}$ can be seen in Figure 6 to correlate with an increase in publishers’ and a decrease in users’ welfare, and indeed when the power parameter $a$ approaches 00, the root proportional ranking function approaches a uniform ranking. That is: $r^{root}_{i}\to\frac{1}{n}$ as $a\to 0$. Similarly, both the intercept parameter $b$ of $r^{lin}$ and the shift parameter $c$ of $r^{log}$ are positively correlated with the publishers’ welfare and negatively correlated with the users’ welfare, and indeed $r^{lin}_{i}\to\frac{1}{n}$ as $b\to\infty$ and $r^{log}_{i}\to\frac{1}{n}$ as $c\to\infty$.
@@ 6 Discussion
We studied an information retrieval game in which publishers have initial documents they prefer to provide and initiated the analysis of no-regret dynamics in such games. We defined the class of proportional ranking functions, which are ranking functions that are determined by a one-variable scalar activation function, and established a full characterization of which proportional ranking functions induce socially-concave games and which induce concave games. We then used this characterization to conclude that any concave activation function guarantees the convergence of no-regret dynamics, thereby offering a system designer a rich family of ranking functions that, under the assumption that publishers minimize their regret, guarantee the reach of the system to a stable state. We then empirically investigated the publishers’ welfare and the users’ welfare, as well as the convergence rate, which measures how long it takes the ecosystem to reach stability. In particular, we examined how these three criteria are affected by the choice of the ranking function and by changes in the ecosystem structure.
@@ A.1 Proof of Lemma 1
$g(x)$ $\coloneqq\sum_{i=1}^{n}\alpha_{i}\cdot u_{i}(x)=\mathbb{E}_{x^{*}\sim P^{*}}\left[\sum_{i=1}^{n}\frac{1}{n}\cdot\left(r_{i}(x;x^{*})-\lambda_{i}\cdot d(x_{i},x_{0}^{i})\right)\right]$ (6)
@@
since $\sum_{i=1}^{n}r_{i}(x;x^{*})=1$ for any strategy profile $x\in X$ and information need $x^{*}\in[0,1]^{k}$. Now, $g(x)$ is concave since for every $i\in N$ and $x_{0}^{i}\in[0,1]^{k}$, $d(x_{i},x_{0}^{i})$ is convex in $x_{i}$. Regarding condition A2., it can be easily shown (using the linearity and monotonicity of expectation) that the convexity of $r_{i}(x;x^{*})$ in $x_{-i}$ for any $x_{i}\in X_{i},x^{*}\in[0,1]^{k}$ implies that $\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. Together with the fact that $d(x_{i},x_{0}^{i})$ is fixed in $x_{-i}$, we get that $u_{i}(x)=\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})-\lambda_{i}\cdot d(x_{i},x_{0}^{i})\right]$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. ∎
A.2 Full derivation of inequality $f^{\prime\prime}(\hat{a})>0$ from the proof of Theorem 1
This appendix provides a detailed derivation of the equation $f^{\prime\prime}(\hat{a})>0$ from the proof of the main result. By the definition of the utility in publishers’ games, and since in the game we constructed the distribution over the information needs is a one-hot distribution that places probability $1$ on $x^{*}=0$:
@@ A.3 Proof that if $d$ is strictly bi-convex and $g$ is concave then $r$ induces strictly concave games
$u_{i}(x_{i},x_{-i})\coloneqq\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]-\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ (18)